Q2: The demand for a single-price monopolist’s product is Q = 60 – 2P where Q is measured in units and P is measured in $/unit.

a) At which price is the demand for the monopolist’s product unit elastic?

b) At which prices is the demand for the monopolist’s product elastic?

c) At which prices is demand for the monopolist’s product inelastic?

d) Suppose the monopoly is currently producing and selling 50 units of output.

  • What price must the monopoly be charging for each unit of output?
  • How does its total revenue, total cost and economic profit change (increase, decrease or unknown) if the monopoly increases its price by $4/unit?
  • Can the monopoly be maximizing economic profit when it is producing and selling 50 units of output?

e) Suppose instead the monopoly is currently producing and selling 24 units of output?

  • What price must the monopoly be charging for each unit of output?
  • How does its total revenue, total cost and economic profit change (increase, decrease or unknown) if the monopoly increases its price by $4/unit?
  • Can the monopoly be maximizing economic profit when it is producing and selling 24 units of output?
Answer #1

question (2) a=60-2P = 2P = 6o-I P= 6o-9 30-0.5 Q 2 TR = Pos (30-0.50 la MR = dTR da 3o-Q =) R = 30-a P, Mc 30 Elastic unit e(a) pemand is unit elastic where MR = 0 Cakula ting for what quantity mp-o Ona Then from Demand Curve corres Corresponding PsBecause for peis The demand is unit elastic (C) For Price p less than or greater Than equals To Zero The demand is inelasticwhether economic Profit increases or decreases That is information about economic Profit is unknown.